OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = 2x - 1 and that pass through the point (4, 1).

OpenStudy (anonymous):

slope of \(y=mx+b\) is \(m\) and therefore the slope of \(y=2x-1\) is \(2\) perpendicular tells you your new slope will be \(-\frac{1}{m}\) in this case \(-\frac{1}{2}\)

OpenStudy (anonymous):

you have the point, it is \((4,1)\) and the slope, it is \(-\frac{1}{2}\) use the "point slope" formula \[y-y_1=m(x-x_1)\] with \[x_1=4,y_1=1,m=-\frac{1}{2}\] to get your answer make a direct substitution

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw